# Efficient Individually Rational Recommender System under Stochastic Order

### v1

@@ 5 Strategic Agents and Bayesian Incentive Compatibility
In this section, we address the case of strategic agents: Agents only see the mechanism’s recommendation, but are free to select any arm. In Subsection 5.1, we lie several observations and assumptions to allow exploration. Then, Subsection 5.2 presents IC-IREGB, which is implemented in Algorithm 3. To simplify our analysis, we shall assume that the rewards are bounded. Namely, there exists $H\in\mathbb{R}$ such that for every $a_{i}\in A^{+}$ it holds that $\left|X(a_{i})\right|\leq H$ almost surely.
@@ 5.1 Enabling Exploration and Harmless Mechanisms
If Assumption 5.1 does not hold for some pair $i,j$, agents would never agree to explore arm $a_{j}$ even though it could hide a higher reward than $a_{i}$.
The main technique used in previous work [25, 26] is hidden exploration: Crafting the mechanism in such a way that agents cannot determine whether they are exploring or exploiting. As a result, a crucial element that affects BIC is the agents’ knowledge about the order of arrival. In the BIC constraint in Definition 2.1, the expectation is conditioned on all the information the agent has about their arrival order. For instance, assume that agents have a uniform belief about their order of arrival and fix any a priori inferior arm $a$. Notice that there is at most one round of exploration for each arm $a$ and potentially many exploitation rounds (under the event that $a$ is observed to be the best arm). Consequently, given enough agents, the probability of exploitation is higher than exploration from an agent’s standpoint.
Proposition 5.1.
@@
Definition 5.1 (Harmless Mechanism).
@@
Observation 5.1.
Assume that $X(a_{0})=0$ almost surely and that $\max_{a_{i}\in\textnormal{pos}(A)}X_{i}<0$. If an IR and BIC mechanism explores the arms of $\textnormal{neg}(A)$, it is necessarily harmful.
We now explain why this observation holds. Fix any harmless mechanism and assume w.l.o.g. that the arms are sorted in decreasing order of expectation. The first agent, knowing she is the first, will only agree to pull $a_{1}$, the a priori optimal arm. Therefore, any IC mechanism will recommend the portfolio $\boldsymbol{p}_{1,1}$. Next, consider the second agent. She knows that the only arm that has been explored so far is $a_{1}$. Thus, if the mechanism recommends anything but $a_{1}$, she will pick $a_{2}$. Since $X(a_{1})<0$ due to the assumptions of the observation ($\max_{a_{i}\in\textnormal{pos}(A)}X_{i}<0$), any harmless mechanism cannot mix $a_{1}$ in a portfolio. Thus, the second agent pulls $a_{2}$. This argument holds inductively for all arms in $\textnormal{pos}(A)$. From here on, we focus on harmless mechanisms only.
In light of Observation 5.1, we need a stronger form of information asymmetry regarding the default arm; hence, we shall make the standard assumption that the default arm $a_{0}$ is a priori superior [5, 23]. This reflects cases where agents are generally informed about the attractiveness of the arms and prefer the default arm over the others.
Assumption 5.2.
The expected value of the rewards satisfy $\mu(a_{0})>\mu(a_{1})\geq\mu(a_{2})\cdots\geq\mu(a_{K})$.
For completeness, we also remark that the IR constraint from Inequality (2) becomes
@@
Notably, the right-hand side is no longer zero. Due to Assumption 5.2, any IR mechanism must recommend the default arm to the first agent.
5.2 The IC-IREGB Algorithm
Input: the IREGB algorithm
1 Initialize an instance of IREGB and update it after every recommendation
2 Recommend the default arm $a_{0}$ to the first agent, observe $x(a_{0})$
3 if $x(a_{0})<\mu_{K}$ then
4      recommend as Greedy to agents $12,\dots,K+$ 1
5
6else
7      recommend $a_{0}$ to agents $12,\dots,K+$ 1
8
9Split the remaining rounds into consecutive phases of $B$ rounds each
10 for phase $k=1,\dots$ do
11      if IREGB exploits (Lines 2-2 in IREGB) then
12            follow IREGB
13
14      else
@@ 2 Recommend the default arm $a_{0}$ to the first agent, observe $x(a_{0})$
15            Pick one agent $l(k)$ from the $B$ agents in this phase uniformly at random and recommend her according to IREGB
16             if an arm $a_{i}$ with $x(a_{i})>x(a_{0})$ was revealed  then
17                  recommend as Greedy
18
19            else
20                  recommend $a_{0}$
21
ALGORITHM 3 Incentive Compatible IREGB (IC-IREGB)
@@
In this subsection, we present a Bayesian incentive-compatible mechanism with approximately optimal welfare. The algorithm, which we call IC-IREGB and is implemented in Algorithm 3, uses IREGB to conduct exploration under the IR and BIC constraints. From a technical standpoint, our approach builds on the techniques introduced by Mansour et al. [25]. We divide rounds into phases: In each phase, there is at most one exploration round (following IREGB). The other rounds are either recommending the default arm $a_{0}$ or greedy exploitation. We denote by Greedy the algorithm that picks the best arm in expectation over the information it has (could be an already observed arm).
@@
The algorithm works as follows. It initializes an IREGB instance in Line 3. For the first agent to arrive, it recommends the default arm and observes its realized reward in Line 3. Then, if the default arm realizes a very low reward (Line 3), we continue by following Greedy for the next $K$ agents (Line 3). Otherwise, in Line 3, we recommend the default arm for those agents. Line 3 splits the next rounds into phases, each with $B$ rounds (we determine $B$ later on). The for loop in Line 3 contains two possible behaviors. If IREGB exploits (the “if” clause in Line 3), we exploit as well. Otherwise, we enter the ‘else” clause in Line 3. One of the agents in this phase, who we note by $l(k)$ and is chosen uniformly at random, will get a recommendation according to IREGB. The other agents will either get the recommendation of Greedy or the default arm. The former refers to the “if” clause in Line 3: As long as we discover an arm $a_{i}$ with $x(a_{i})>x(a_{0})$, which means that IREGB can explore all arms using Bernoulli trails. This completes the description of the algorithm.
@@
To analyze the social welfare of IC-IREGB, we need to determine the phase length $B$; thus, we introduce the following quantities $\xi$ and $\gamma$. Assumption 5.1 hints that there exist $\xi>0$ and $\gamma>0$ such that every arm $i\in A^{+}$ has a chance of at least $\gamma$ to be greater than any other arm by at least $\xi$. Formally, for all $i\in A^{+}$, it holds that $\Pr(\forall i^{\prime}\in A^{+}\setminus\{i\}:\mu_{i}-X_{i^{\prime}}>\xi)>\gamma$. We use these quantities and set $B=\left\lceil\frac{H}{\xi\gamma}\right\rceil+1$. We are ready to summarize the properties of IC-IREGB.
Theorem 5.1.
Under Assumptions 2.1,5.1 and 5.2, IC-IREGB satisfies IR and BIC. In addition,
@@
$SW_{T}({\small\textnormal{{IC-IREGB}}})\geq\left(1-O\left(\frac{K\eta H\operatorname{\mathbb{E}}\left[\frac{1}{\delta}\right]}{T\xi\gamma}\right)\right){\small\textnormal{{OPT}}}_{\infty}.$
@@ 6 Conclusion and Discussion
Our model emphasizes the individual rationality requirement, ensuring that agents receive in expectation at least as much as they would have received without using the mechanism, based on the mechanism’s knowledge. In Section 3, we introduced an auxiliary GMDP in Section 3 and achieved our main contribution: An index-based optimal policy that only requires sorting arms according to their expected values. Later, in Section 4, we leveraged this optimal GMDP policy to develop an approximately optimal algorithm and analyzed its convergence rate. Finally, in Section 5, we addressed incentive compatibility.
@@
We see considerable scope for future work. On the technical side, one possible follow-up is to relax Assumption 2.1; more details can be found in Subsection 6.1. Another possible direction is to consider stochastic bandits with Bayesian priors, moving beyond the assumption of static rewards. Our individual rationality constraint is readily extendable to this case as well. Conceptually, our definition of individual rationality concerns the expected value. However, in many settings, other factors, such as variance, should also be considered. In such cases, our constraint can be generalized to $\sum_{a_{i}\in A}\boldsymbol{p}(a_{i})\operatorname{\mathbb{E}}\left[f(X(a_{i}))\mid\mathcal{I}\right]\geq 0$, for some function $f$. This more general formulation can express risk-aversion or risk-seeking behavior, as well as more complex quantities that depend on the reward distribution. Our results from Section 3 hold for any general $f$, as long as it agrees with the stochastic order. Finally, our IR notion could be applied to other explore-exploit models, such as Markov Decision Processes.
6.1 Open Problem: Relaxing Assumption 2.1
@@
Unfortunately, our results hold only under the stochastic order assumption, Assumption 2.1. This limitation is not merely a byproduct of our proof technique; there are examples where OGP ceases to be optimal due to incorrect planning (see the proof of Proposition F.3 in the appendix).
@@
To illustrate why OGP leads to sub-optimal performance in the absence of stochastic order, consider three crucial aspects of selecting an arm $a_{j}$ as part of a portfolio (a randomized action). First, its expected value $\mu(a_{j})$ determines the likelihood of exploring it. Second, the probability of obtaining a positive value, $\Pr(X_{a_{j}}>0)$, influences the exploration of all other arms via Bernoulli trials. Third, the potential for exceptionally high rewards if $X_{a_{j}}$ turns out to be positive, $\operatorname{\mathbb{E}}(\max_{a_{l}}X_{a_{l}}\mid X_{a_{j}}>0)$, can generate significant regret if not explored. Under Assumption 2.1, ordering the arms in $\textnormal{neg}(A)$ by each of these factors yields the same order. Without stochastic order, the exploration index, and whether such an index even exists, is unclear. We conjecture that a different index-like policy is optimal.
Conjecture 6.1.
@@
Let $A$ be an arbitrary set of arms, not necessarily satisfying Assumption 2.1. For every state $U\subseteq A$, let $f^{U}$ be the real-valued function, $f^{U}:\textnormal{neg}(U)\rightarrow\mathbb{R}$ defined as
@@ Acknowledgments
The work of O. Ben-Porat was supported by the Israel Science Foundation (ISF; Grant No. 3079/24).
Appendix A Proof Outline for Theorem 3.1
@@
A.1 Preliminaries
We denote the set of all states by $\mathcal{S}=2^{A}$. A policy is a mapping from previous states and actions to a randomized action. Formally, let $\mathcal{H}$ be the set of histories, $\mathcal{H}=\cup_{k=0}^{K}\left(\mathcal{S}\times\Delta(A)\right)^{k}$ be a tuple comprising pairs of states and randomized actions taken. A policy $\pi$ is a function $\pi:\mathcal{H}\times\mathcal{S}\rightarrow\Delta(A)$. We say that a policy is IR if for every $h\in\mathcal{H},s\in\mathcal{S}$, $\pi(h,s)\in\textit{IR}(s)$. From here on, we consider IR policies solely. Given a policy $\pi$ and a pair $(h,s)$, we let $W(\pi,h,s)$ denote the expected reward of $\pi$ when starting from $s$ after witnessing $h$. Namely,
$W(\pi,h,s)=\begin{cases}R(s)&\textnormal{if }\textit{IR}(s)=\emptyset\\ \sum_{a\in s}\pi(h,s)(a)W(\pi,h\oplus(s,a),s\setminus\{a\})&\textnormal{otherwise}\end{cases}.$ (7)
For every state $s$, let $W^{\star}(s)=\sup_{\pi^{\prime}}W(\pi^{\prime},\emptyset,s)$. [^6] As we show in Proposition A.1, there exists a policy that attains this supremum. While policies may depend on histories, it often suffices to consider stationary policies.
Definition A.1 (Stationary).
An IR policy $\pi$ is stationary if for every two histories $h,h^{\prime}\in\mathcal{H}$ and a state $s\in\mathcal{S}$, $\pi(h,s)=\pi(h^{\prime},s)$.
@@ A.1 Preliminaries
A.2 Binary Structure
A structural property of the above GMDP is that in every terminal state $s_{t}$, $\textnormal{pos}(s_{t})=\emptyset$, or otherwise we could explore more arms; thus, intuitively, the arms in $\textnormal{pos}(A)$ provide us “power” to explore the arms of $\textnormal{neg}(A)$. Following this logic, in every state we should aim to explore arms from $\textnormal{neg}(A)$ and not those of $\textnormal{pos}(A)$, subject to satisfying the IR constraint.

### v2

@@ 5 Strategic Agents and Bayesian Incentive Compatibility
In this section, we address the case of strategic agents: Agents only observe the mechanism’s recommendation, but are free to select any arm. In Subsection 5.1, we lay out several observations and assumptions that enable exploration. Then, Subsection 5.2 presents BIC-IREGB, which is implemented in Algorithm 2. To simplify our analysis, we shall assume that the rewards are bounded. Namely, there exists $H\in\mathbb{R}$ such that for every $a_{i}\in A^{+}$, it holds that $\left|X(a_{i})\right|\leq H$ almost surely.
@@ 5.1 Enabling Exploration and Harmless Mechanisms
If Assumption 2 does not hold for some pair $i,j$, agents would never agree to explore arm $a_{j}$ even though it could hide a higher reward than $a_{i}$.
The main technique used in previous work [25, 26] is hidden exploration: Crafting the mechanism in such a way that agents cannot determine whether they are exploring or exploiting. As a result, a crucial element that affects BIC is the agents’ knowledge about the order of arrival. In the BIC constraint in Definition 1, the expectation is conditioned on all the information the agent has about their arrival order. For instance, assume that agents have a uniform belief about their order of arrival and fix any a priori inferior arm $a$. Notice that there is at most one round of exploration for each arm $a$ and potentially many exploitation rounds (under the event that $a$ is observed to be the best arm). Consequently, given enough agents, the probability of exploitation is higher than exploration from an agent’s standpoint.
Proposition 4
@@
Definition 4 (Harmless Mechanism)
@@
Observation 4
Assume that $X(a_{0})=0$ almost surely and that $\max_{a_{i}\in\textnormal{pos}(A)}X_{i}<0$. If a MIR and BIC mechanism explores the arms of $\textnormal{neg}(A)$, it is necessarily harmful.
We now explain why this observation holds. Fix any harmless mechanism and assume w.l.o.g. that the arms are sorted in decreasing order of expectation. The first agent, knowing she is the first, will only agree to pull $a_{1}$, the a priori optimal arm. Therefore, any BIC mechanism will recommend the portfolio $\boldsymbol{p}_{1,1}$. Next, consider the second agent. She knows that the only arm that has been explored so far is $a_{1}$. Thus, if the mechanism recommends anything but $a_{1}$, she will pick $a_{2}$. Since $X(a_{1})<0$ due to the assumptions of the observation ($\max_{a_{i}\in\textnormal{pos}(A)}X_{i}<0$), any harmless mechanism cannot mix $a_{1}$ in a portfolio. Thus, the second agent pulls $a_{2}$. This argument holds inductively for all arms in $\textnormal{pos}(A)$. From here on, we focus on harmless mechanisms only.
In light of Observation 4, we need a stronger form of information asymmetry regarding the default arm; hence, we shall make the standard assumption that the default arm $a_{0}$ is a priori superior [6, 23]. This reflects cases where agents are generally informed about the attractiveness of the arms and prefer the default arm over the others.
Assumption 3
The expected values of the rewards satisfy $\mu(a_{0})>\mu(a_{1})\geq\mu(a_{2})\cdots\geq\mu(a_{K})$.
For completeness, we also remark that the MIR constraint from Inequality (2) becomes
@@
Notably, the right-hand side is no longer zero. Due to Assumption 3, any MIR mechanism must recommend the default arm to the first agent.
5.2 The BIC-IREGB Algorithm
Algorithm 2 Incentive Compatible IREGB (BIC-IREGB)
0:  the IREGB algorithm
1:  initialize an instance of IREGB and update it after every recommendation
2:  recommend the default arm $a_{0}$ to the first agent, observe $x(a_{0})$
3:  if $x(a_{0})<\mu_{K}$  then
4:     recommend using Greedy to agents $12,\dots,K+$ 1
5:  else
6:     recommend $a_{0}$ to agents $12,\dots,K+$ 1
7:  split the remaining rounds into consecutive phases of $B$ rounds each
8:  for phase $k=1,\dots$  do
9:     if IREGB exploits (Lines 8-9 in IREGB)  then
10:        follow IREGB
11:     else
@@ 5.2 The BIC-IREGB Algorithm
12:        pick one agent $l(k)$ from the $B$ agents in this phase uniformly at random and recommend her according to IREGB
13:        as for the rest of the agents,
14:        if an arm $a_{i}$ with $x(a_{i})>x(a_{0})$ was revealed  then
15:           recommend using Greedy
16:        else
17:           recommend $a_{0}$
@@
In this subsection, we present a Bayesian incentive-compatible mechanism with approximately optimal welfare. The algorithm, which we call BIC-IREGB and is implemented in Algorithm 2, uses IREGB to conduct exploration under the MIR and BIC constraints. From a technical standpoint, our approach builds on the techniques introduced by Mansour et al. [25]. We divide rounds into phases: In each phase, there is at most one exploration round (following IREGB). The other rounds are either recommending the default arm $a_{0}$ or greedy exploitation. We denote by Greedy the algorithm that picks the best arm in expectation over the information it has (could be an already observed arm).
@@
The algorithm works as follows. It initializes an IREGB instance in Line 1. For the first agent to arrive, it recommends the default arm and observes its realized reward in Line 2. Then, if the default arm realizes a very low reward, we enter the “if” clause in Line 3 and continue by following Greedy for the next $K$ agents (Line 4). Otherwise, in Line 6, we recommend the default arm for those agents. Line 7 splits the next rounds into phases, each with $B$ rounds (we determine $B$ later on). The for loop in Line 8 contains two possible behaviors. If IREGB exploits (the “if” clause in Line 9), we exploit as well. Otherwise, we enter the “else” clause in Line 11. One of the agents in this phase, who we note by $l(k)$ and is chosen uniformly at random, will get a recommendation according to IREGB (Line 12). The other agents will either get the recommendation of Greedy or the default arm. The former refers to the “if” clause in Line 14: As long as we discover an arm $a_{i}$ with $x(a_{i})>x(a_{0})$, which means that IREGB can explore all arms using Bernoulli trials. This completes the description of the algorithm.
@@
To analyze the social welfare of BIC-IREGB, we need to determine the phase length $B$; thus, we introduce the following quantities $\xi$ and $\gamma$. Assumption 2 hints that there exist $\xi>0$ and $\gamma>0$ such that every arm $i\in A^{+}$ has a chance of at least $\gamma$ to be greater than any other arm by at least $\xi$. Formally, for all $i\in A^{+}$, it holds that $\Pr(\forall i^{\prime}\in A^{+}\setminus\{i\}:\mu_{i}-X_{i^{\prime}}>\xi)>\gamma$. We use these quantities and set $B=\left\lceil\frac{H}{\xi\gamma}\right\rceil+1$. We are ready to summarize the properties of BIC-IREGB.
Theorem 5.1
Under Assumptions 1,2 and 3, BIC-IREGB satisfies MIR and BIC. In addition,
@@
$SW_{T}({\small\textnormal{{BIC-IREGB}}})\geq\left(1-O\left(\frac{K\eta H\operatorname{\mathbb{E}}\left[\frac{1}{\delta}\right]}{T\xi\gamma}\right)\right){\small\textnormal{{OPT}}}_{\infty}.$
@@ 6 Conclusion and Discussion
Our model emphasizes the MIR requirement, ensuring that agents receive, in expectation, at least as much as they would have received without using the mechanism, based on the mechanism’s accumulated information. In Section 3, we introduced an auxiliary GMDP and achieved our main contribution: An index-based optimal policy that only requires sorting arms according to their expected values. Later, in Section 4, we leveraged this optimal GMDP policy to develop an approximately optimal algorithm and analyzed its convergence rate. Finally, in Section 5, we addressed incentive compatibility.
@@
We see considerable scope for future work. On the technical side, one possible follow-up is to relax Assumption 1; more details can be found in Subsection 6.1. Another possible direction is to consider stochastic bandits with Bayesian priors, moving beyond the assumption of static rewards. The MIR constraint is readily extendable to this case as well. More conceptually, MIR concerns only the expected value. However, in many settings, other factors, such as variance, should also be considered. In such cases, our constraint can be generalized to $\sum_{a_{i}\in A}\boldsymbol{p}(a_{i})\operatorname{\mathbb{E}}\left[f(X(a_{i}))\mid\mathcal{I}\right]\geq 0$, for some function $f$. This more general formulation can express risk-aversion or risk-seeking behavior, as well as more complex quantities that depend on the reward distribution. Our results from Section 3 hold for any general $f$, as long as it agrees with the stochastic order. Finally, the MIR constraint could be applied to other explore-exploit models, such as Markov Decision Processes.
6.1 Open Problem: Relaxing Assumption 1
@@
Unfortunately, our results hold only under the stochastic order assumption, Assumption 1. This limitation is not merely a byproduct of our proof technique; there are examples where OGP ceases to be optimal due to incorrect planning (see the proof of Proposition 15 in the appendix).
@@
To illustrate why OGP leads to sub-optimal performance in the absence of stochastic order, consider three crucial aspects of selecting an arm $a_{j}$ as part of a portfolio (a randomized action). First, its expected value $\mu(a_{j})$ determines the likelihood of exploring it. Second, the probability of obtaining a positive value, $\Pr(X_{a_{j}}>0)$, influences the exploration of all other arms via Bernoulli trials. Third, the potential for exceptionally high rewards if $X_{a_{j}}$ turns out to be positive, $\operatorname{\mathbb{E}}(\max_{a_{l}}X_{a_{l}}\mid X_{a_{j}}>0)$, can generate significant regret if not explored. Under Assumption 1, ordering the arms in $\textnormal{neg}(A)$ by each of these factors yields the same order. Without stochastic order, the exploration index, and whether such an index even exists, is unclear. We conjecture that a different index-like policy is optimal.
Conjecture 1
@@
Let $A$ be an arbitrary set of arms, not necessarily satisfying Assumption 1. For every state $U\subseteq A$, let $f^{U}$ be the real-valued function, $f^{U}:\textnormal{neg}(U)\rightarrow\mathbb{R}$ defined as
@@ Acknowledgments
The work of Omer Ben-Porat was supported by the Israel Science Foundation (ISF; Grant No. 3079/24).
Appendix 0.A Proof Outline for Theorem 3.1
@@
0.A.1 Preliminaries
We denote the set of all states by $\mathcal{S}=2^{A}$. A policy is a mapping from previous states and actions to a randomized action. Formally, let $\mathcal{H}$ be the set of histories, $\mathcal{H}=\cup_{k=0}^{K}\left(\mathcal{S}\times\Delta(A)\right)^{k}$ be a tuple comprising pairs of states and randomized actions taken. A policy $\pi$ is a function $\pi:\mathcal{H}\times\mathcal{S}\rightarrow\Delta(A)$. We say that a policy is MIR if for every $h\in\mathcal{H},s\in\mathcal{S}$, $\pi(h,s)\in\textit{MIR}(s)$. From here on, we consider MIR policies solely. Given a policy $\pi$ and a pair $(h,s)$, we let $W(\pi,h,s)$ denote the expected reward of $\pi$ when starting from $s$ after witnessing $h$. Namely,
$W(\pi,h,s)=\begin{cases}R(s)&\textnormal{if }\textit{MIR}(s)=\emptyset\\ \sum_{a\in s}\pi(h,s)(a)W(\pi,h\oplus(s,a),s\setminus\{a\})&\textnormal{otherwise}\end{cases}.$ (7)
For every state $s$, let $W^{\star}(s)=\sup_{\pi^{\prime}}W(\pi^{\prime},\emptyset,s)$. [^6] As we show in Proposition 5, there exists a policy that attains this supremum. While policies may depend on histories, it often suffices to consider stationary policies.
Definition 5 (Stationary)
A MIR policy $\pi$ is stationary if for every two histories $h,h^{\prime}\in\mathcal{H}$ and a state $s\in\mathcal{S}$, $\pi(h,s)=\pi(h^{\prime},s)$.
@@ 0.A.1 Preliminaries
0.A.2 Binary Structure
A structural property of the above GMDP is that in every terminal state $s_{t}$, $\textnormal{pos}(s_{t})=\emptyset$, or otherwise we could explore more arms; thus, intuitively, the arms in $\textnormal{pos}(A)$ provide us “power” to explore the arms of $\textnormal{neg}(A)$. Following this logic, in every state we should aim to explore arms from $\textnormal{neg}(A)$ and not those of $\textnormal{pos}(A)$, subject to satisfying the MIR constraint.